OpenStudy (anonymous):

josh and mae work at a concession stand. thy each earn $8 per hour. josh worked 3 hours more than mae. if josh and mae earned a total of $120 , how many hours did josh work ???

OpenStudy (anonymous):

8(t+3)+8(t)=120

OpenStudy (anonymous):

thank youu :-)

OpenStudy (anonymous):

how do ii*figure out T ?

OpenStudy (anonymous):

Distributing above equation 8t+24+8t=120 16t+24=120 16t=120-24 16t=96 t=6

OpenStudy (anonymous):

ooohhh ooh kay good looks